OpenStudy (anonymous):

What is the area of a regular hexagon with a side length of 5.2 centimeters and a distance from the center to a vertex of 5.2 centimeters?

OpenStudy (anonymous):

area of 1 equilateral = (5.2)^2 (sqrt3)/4 area of the hexagon = 6(2.6)^2 ( sqrt3)